OpenStudy (anonymous):

What is the magnitude of the following complex number: (11+67i)/(1+3i)

OpenStudy (anonymous):

\[\left| \frac {11+67i}{1+3i} \right|=\left| \frac {11+67i}{1+3i}\frac{1-3i}{1-3i} \right|=\left| \frac{(11+201)+(67-33)i}{10} \right|\]\[=\left|\frac{106+17i}{5} \right|=\sqrt{\frac{106^2+17^2}{25}} approx 4.29 \]

OpenStudy (anonymous):

I hate calculators... real answer is approx 21.5

OpenStudy (anonymous):

yeah thats the one, Thanks!

OpenStudy (anonymous):

No sweat. Do math every day.